Grade 99 grade science

A drunkard is walking along a straight road. He takes five steps forward and three steps backward and so
on. Each step is 1 m long and takes 1 s. There is a pit on the road 11 m away from the starting point. The
drunkard will fall into the pit after

Vaibhav Maheshwari
4 Years agoGrade 9

1 Answer

Mayank Ranka
4 Years ago
He moves 5 steps forward and 3 steps backward in 8sec after that he is 2m away from an initial position
x= 5-3=2m in 8 sec
after 24sec he will 6m away from an initial position
He moves further 5m forward then he falls in a pit which is 11m away from the initial position.
total time T=24+5=29sec.